OpenStudy (anonymous):

guys plz tel me how to calculate the number which comes before roots i mean like this 8√4

OpenStudy (saifoo.khan):

\[8\sqrt4\]\[8 \times 2\]\[16\]

OpenStudy (anonymous):

i am going to bet that the question was \[\sqrt[8]{4}\] in which case the answer would be \[\sqrt[4]{2}\] in simplest radical form

OpenStudy (anonymous):

that is because you have \[\sqrt[8]{2^2}\] and so you can divide the "index" which is 8, by the exponent, which is two.

OpenStudy (anonymous):

dude answer this 1-10√5000/105000 ?

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

\[1- \sqrt[8]{5000/105000}\] answer now